How much of a 60% acid solution must be mixed with 50 gallons of a 24% acid solution to obtain a solution that is 50% acid?
Answer by lwsshak3(11628)

How much of a 60% acid solution must be mixed with 50 gallons of a 24% acid solution to obtain a solution that is 50% acid?
***
let x=amt of 60% solution to mix
50+x=amt of final mixture
..
60%x+24%*50=50%(50+x)
.60x+12=25+.50x
.10x=13
x=130
amt of 60% solution to mix=130 gal